Circular No. 7724

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001en IN NGC 523 D. Hutchings and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Sept. 26.5 (mag ∼ 17.5) and 27.3 UT (mag ∼ 17.0). SN 2001en is located at α = 1^{h}25^{m}22^{s}.90, δ = +34^{o}01′30′′.5 (equinox 2000.0), which is 6′′.4 east and 2′′.8 south of the eastern nucleus of NGC 523. An unfiltered KAIT image taken on Sept. 20.4 showed nothing at this positon (limiting mag ∼ 19.5).

SUPERNOVA 2001el IN NGC 1448 L. Wang and D. N. Kasen, Lawrence Berkeley National Laboratory

(LBNL); D. Baade, European Southern Observatory (ESO); C. Fransson, Stockholm Observatory; D. A. Howell, LBNL; P. Hoeflich, University of Texas; P. Lundqvist, Stockholm Observatory; P. E. Nugent, LBNL; and J. C. Wheeler, University of Texas, communicate: “High-quality spectropolarimetry (range 417–860 nm; spectral resolution 1.27 nm and 0.265 nm/pixel) of SN 2001el was obtained with the ESO Very Large Telescope UT3 (+ FORS1) on Sept. 26 UT. The spectra confirm (cf. IAUC 7723) that this is a type-Ia supernova before maximum, with a normal spectrum similar to that of SN 1994D except for a strong double-troughed absorption feature around 800.0 nm (FWHM ∼ 22 nm). The absorption dips of Si II (635.5, 564.0 nm) and Fe II (492.4, 516.9 nm) all show velocities of ∼ 10 000 km/s. The supernova was polarized at a level of 0.9 percent (uncorrected for interstellar polarization), with clearly polarized features across several of the spectral lines. Ca II lines are present in the spectrum at a velocity comparable to that of the Si II 635.5-nm feature but are much weaker than the absorption feature at 800.0 nm. It is possible that the 800-nm feature is a second component of high-velocity Ca II, as suggested by Hatano et al. (1999, Ap.J. 525, 881) to explain a similar (but much weaker) feature in SN 1994D. If identified with Ca II, the feature would be due to a detatched shell/clump with a velocity range from 20 000 to 26 000 km/s. The 800.0-nm feature also shows a significantly higher degree of polarization than other spectral lines.”

2001 September 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green